# Using Large Language Models to Tackle Fundamental Challenges in Graph Learning: A Comprehensive Survey

### v1

@@ 3.4.4 Summary of Dynamic Instability
4 Future Directions and Open Challenges
4.1 Incompleteness in Graphs
4.1.1 Controlling Hallucination in LLM-based Graph Completion
LLMs hold strong potential for completing missing nodes, edges, or attributes in graphs by generating latent knowledge from textual context. However, a key challenge lies in mitigating hallucination—the generation of plausible but incorrect or structurally inconsistent information. Ensuring that LLM-generated content aligns with the graph’s schema and topology requires future work on structure-constrained decoding, confidence-aware integration, and post-generation validation using GNNs or external knowledge bases. Jointly training LLMs with graph-aware objectives may further reduce hallucination and improve the structural fidelity of completions.
4.1.2 LLMs’ Robust Inference on Incomplete Data
LLMs demonstrate robustness when faced with missing or noisy inputs, but their reasoning remains opaque and uncontrolled. Future research should investigate how to guide LLMs toward faithful, graph-consistent inference, particularly when data is partially observed. One avenue is to design structure-aware prompts or adapters that encode local subgraph structure, enabling LLMs to learn inductive biases from graph topology rather than relying solely on pretraining knowledge. Additionally, LLMs could be enhanced to support iterative reasoning over subgraphs, where initial predictions are refined through multiple interaction rounds with the graph context. This would mimic how GNNs aggregate information layer by layer, and could bridge the reasoning styles of symbolic and connectionist methods.
4.1.3 Benchmarks and Interpretability in LLM-based Graph Learning
Despite growing interest, there is a lack of standardized benchmarks and evaluation protocols tailored for LLM-based graph learning. Future work should establish datasets and tasks that reflect real-world challenges such as incomplete graphs, noisy node attributes, and open-world reasoning. Furthermore, interpretability remains a major hurdle: how can we trust LLM-driven graph predictions? Tools that visualize attention over graph structures, track reasoning chains within LLMs, or contrast outputs under counterfactual graph edits could greatly enhance model transparency.
4.2 Imbalance in Graphs
4.2.1 Semantic-Topology Joint Debias
To address the imbalance between semantic logic and topological patterns in graph data, the focus of research is on constructing a text-graph joint representation space to uncover the deep interconnections between multimodal information. Leveraging the causal reasoning ability of LLMs, it is possible to identify the semantic roots of structural imbalance, such as topological sparsity caused by the logical absence of long-tail relationships in knowledge graphs. By generating logically constrained virtual edges or subgraph structures, this approach can reconstruct the global distribution of the graph from a causal perspective, achieving a synergistic enhancement of semantic coherence and topological integrity. Typical application scenarios include: in knowledge graph completion tasks, LLMs can be used to parse the implicit semantic relationships between entities and generate virtual triples; in social network analysis, user behavior text can be combined to infer potential social links, correcting network structural flaws caused by data collection biases.
4.2.2 Dynamic Cross-Modal Transfer
To mitigate class or structural bias in graph learning tasks caused by data sparsity, lightweight adaptation techniques need to be developed for the dynamic transfer of LLMs’ open-domain knowledge to graph structures. By designing cross-modal projection layers and attention alignment mechanisms, LLMs’ semantic representations can be mapped to a unified vector space with graph topological features. In molecular property prediction scenarios, this can achieve spatial matching between atomic-level text descriptions and molecular graphs, or in social network analysis, it can integrate user profile text with community structure features. This technology breaks through the closed training approach of traditional graph models, utilizing LLMs’ generalized knowledge to compensate for deficiencies in domain-specific data.
4.2.3 Trustworthy Generation & Efficiency Optimization
To ensure semantic validity when LLMs generate minority-class nodes or edges in imbalanced graphs, future work should focus on suppressing hallucinations through structure-aware verification. By cross-validating GNN-derived topological patterns with LLM-generated semantics, models can dynamically filter or penalize low-confidence, type-inconsistent outputs. Prototype alignment and discriminative validators further help constrain generation within class-specific semantics. Meanwhile, parameter-efficient tuning (e.g., LoRA) and multimodal pretraining enable scalable deployment and real-time adaptation to long-tail patterns in dynamic graphs, such as correcting abnormal transaction structures via time-aligned textual streams.
4.3 Cross-Domain Heterogeneity in Graphs
4.3.1 Context-Aware Heterogeneous Attribute Learning
Existing LLM-based methods for handling attribute heterogeneity typically treat LLMs or MLLMs as powerful preprocessing tools for generating unified embeddings for node attributes. However, modeling attributes in isolation without leveraging the abundant contextual information leads to suboptimal representations. For instance, the semantic meaning of a textual node attribute can be enriched by incorporating other textual attributes or even multimodal attributes associated with the same node. Additionally, the connectivity patterns of a node offer important relational information that enhance the interpretation of its attributes. Moreover, different tasks emphasize different aspects of attributes, making task-related context essential for generating embeddings aligned with specific learning objectives. Future research can explore integrating semantic, structural, and task-related context information to enhance heterogeneous attribute learning.
4.3.2 Symmetry-Preserving Heterogeneous Structure Learning
LLMs exhibit strong capabilities in understanding and reasoning over heterogeneous graph structures represented in natural language. However, they do not explicitly incorporate graph symmetries such as invariance. For instance, the representation of a graph should remain consistent regardless of the ordering of its input node sequence, as different orderings do not alter the underlying structure. Existing LLM-based methods fail to fully exploit such graph symmetry, making it difficult to determine whether variations in the input graph descriptions stem from distinct heterogeneous structures from different source domains or merely the same structure under different permutation actions. To address this issue, one possible approach is to design prompts that explicitly describe the relevant symmetries and provide examples in the context, helping LLMs understand the symmetric constraints of the input graph. Additionally, LLMs can be fine-tuned using contrastive learning to distinguish whether two differently ordered node sequences represent the same underlying graph structure to capture the symmetry knowledge.
4.3.3 Effective Integration of GNNs and LLMs
Although LLMs excel at capturing semantic and structural knowledge from graph data represented in text, they cannot utilize the entire graph data due to limited context window sizes. In contrast, GNNs can be easily trained on large graphs due to the efficient message passing mechanism. Current methods for handling cross-domain heterogeneity explore using LLMs as attribute encoders for GNNs or generating graph tokens with GNNs for LLM reasoning. These integration methods are limited by the expressiveness of GNNs, which may fail to detect some critical substructures [265, 266]. To better utilize the advantages of both methods, one potential direction is to design a dual architecture, where the final prediction could be based on the outputs of both LLMs and GNNs. Another direction is to distill the knowledge from a GNN trained on the entire graph into an LLM, allowing the LLM to effectively capture the knowledge from the graph data without finetuning on the whole graph.
4.4 Dynamic Instability in Graphs
4.4.1 Querying and Reasoning
A key challenge is scaling LLM-based reasoning to large and high-frequency dynamic graphs while keeping inferences grounded in the underlying graph structure. Current models often struggle with hallucinations and fail to reflect the evolving nature of real-world graphs. Another limitation is the difficulty in representing complex temporal dependencies in a form that LLMs can effectively process. Future work should focus on building tighter integrations between LLMs and GNNs, developing pretraining strategies that incorporate temporal data, and designing efficient, time-aware prompting and retrieval methods. Improving the ability of LLMs to perform causal reasoning and making their decision processes more transparent are also important directions for research.
4.4.2 Generating and Updating
Current LLMs face challenges in generating graph elements that are faithful to known facts, structurally consistent, and aligned with domain constraints. Frequent updates to dynamic graphs also raise concerns about computational cost and system stability. Controlling the generation process remains challenging, especially when outputs need to be incorporated into existing graph structures. Future research should investigate constrained generation techniques, efficient methods for incremental updates, and mechanisms for integrating feedback from downstream tasks. Extending LLMs to model more complex graph dynamics and to incorporate signals from multiple modalities may enhance their ability to support real-world applications.
4.4.3 Evaluation and Application
LLM-based evaluation methods must address issues of reliability, bias, and limited generalization. Static benchmarks fail to reflect the changing nature of dynamic graphs and the models that operate on them. The high cost and latency of LLMs also present barriers to real-time deployment. Future directions include developing adaptive evaluation frameworks that combine LLM-based reasoning with traditional performance metrics and exploring human-in-the-loop strategies to improve evaluation quality. Research should also focus on building lightweight deployment architectures and designing LLM-based systems that adapt proactively to changing graph environments while carefully addressing fairness and ethical concerns.
@@ 5 Conclusion
In this survey, we present a comprehensive review of how recent advances in large language models can address four fundamental challenges in graph learning that arise from mismatches between ideal assumptions and real-world graphs. These core challenges are (1) Incompleteness of graph structures or attributes, (2) severe Imbalance in node/edge distributions, (3) Cross-Domain Heterogeneity in semantics and structure, and (4) Dynamic Instability due to evolving topologies and interactions. Each of these issues highlights a facet of graph data that can confound traditional graph learning methods and limit their generalization ability.
We reviewed a wide array of LLM-integrated approaches proposed to tackle these challenges, explaining them through a unifying conceptual framework. The findings across the literature demonstrate that LLMs offer significant advantages when combined with graph learning techniques. For incomplete graphs, LLMs provide powerful semantic reasoning and external knowledge to infer missing node attributes and relationships, effectively acting as intelligent imputers to complete graphs beyond what purely graph-based methods can do. In imbalanced graphs, LLMs can generate synthetic data and augment feature spaces to alleviate class or structural imbalance, introducing new examples and nuanced features that help models learn minority patterns and thus improve robustness against bias. For heterogeneous graphs, LLMs produce unified semantic embeddings and serve as a bridge across modalities and domains, harmonizing disparate attribute types and structural schemas into a common representation. In dynamic graphs, LLMs enable advanced temporal reasoning and contextual analysis, from anticipating future links to explaining changes, thereby helping graph models continuously adapt over time.
@@
Looking ahead, several open questions and promising directions emerge. First, improving the efficiency of LLM-graph integration remains critical, particularly for large-scale or real-time applications. Second, enhancing the explainability of LLM-driven graph reasoning will foster trust in high-stakes domains like healthcare and finance. Third, developing frameworks for continuous adaptation to dynamic graphs—without catastrophic forgetting—requires further exploration. Fourth, bridging the gap between textual knowledge encoded in LLMs and structural graph signals could unlock novel cross-modal learning approaches. Finally, addressing the issue of hallucination in LLM-based graph generation is essential to ensure that generated content is both semantically accurate and structurally consistent, fostering more reliable and trustworthy graph learning.
@@
In closing, the integration of large language models with graph learning represents a new frontier that brings together the best of two worlds: the deep semantic understanding from NLP and the structured relational modeling from graph ML. The synergy between these fields is already yielding models that are more robust, knowledgeable, and adaptable than those based on either technology alone. As research continues along the directions identified, we anticipate rapid progress toward graph learning systems with far greater generality and intelligence, capable of reasoning about the rich, dynamic networks that underlie real-world data.
@@ Acknowledgements
This work was supported in part by the Shenzhen Science and Technology Program under Grant JCYJ20240813151129038; in part by the National Natural Science Foundation of China under Grant 52172350, and Grant 51775565; in part by the Guangdong Basic and Applied Research Foundation under Grant 2022B1515120072; and in part by the Science and Technology Planning Project of Guangdong Province under Grant 2023B1212060029.

### v2

@@ 3.4.5 Summary of Dynamic Instability
4 Limitations, Challenges, and Future Directions
This section builds upon the preceding analysis to consolidate key challenges and limitations in graph-based learning with LLMs, and to outline potential directions for future research. The discussion is organized around several considerations, including efficiency and scalability, explainability, fairness, and faithfulness, and security, robustness, and governance, which together reflect both the current barriers and the broader goals for advancing LLM–graph integration.
4.1 Efficiency and Scalability
Scalability is a key challenge in integrating LLMs with graph data, given the high memory and computation costs of large, heterogeneous, and dynamic graphs. We summarize efficiency challenges across scenarios and explore ways to improve tractability, compactness, and inference speed.
4.1.1 Incompleteness
LLM-based methods for graph completion show promise in generating missing nodes, edges, or attributes, but they also introduce risks of hallucination and structural inconsistency. To improve the fidelity of generated content, future research should explore structure-constrained decoding, confidence-aware integration, and post-hoc validation using GNNs or external knowledge bases [121]. Incorporating structural priors from graphs directly into training objectives may further enhance the alignment between textual reasoning and graph topology [136]. At the same time, reducing representational redundancy is critical for scalability, motivating the development of compact graph–text co-representations. For example, injecting structure tokens or refactoring sequence layouts can reduce the need for full-graph serialization during inference [366, 367]. Besides, efficiency-oriented designs should also consider concrete resource metrics such as sequence length, memory usage, and per-sample or per-token computation cost [368].
4.1.2 Imbalance
Imbalanced distributions are common in real-world graphs, where a few dominant classes overshadow long-tail entities that often lack sufficient structural and semantic support [188, 187]. Uniformly applying LLMs across all samples is inefficient and may overfit majority patterns while underperforming on minority nodes. Moreover, hallucinations in underrepresented regions pose reliability risks, especially when topological signals are weak [191]. To mitigate these issues, future research should consider selective LLM invocation for hard or minority-class cases, while using GNNs for routine instances and transferring LLM knowledge into lightweight models via parameter-efficient tuning [369, 370]. Structure-aware validation modules that cross-check LLM outputs against graph-derived patterns can improve robustness [133], and a promising direction is to construct joint text–graph causal representations to help models infer the semantic origins of structural sparsity and synthesize logically grounded virtual connections [190].
4.1.3 Cross-Domain Heterogeneity
Graphs that span multiple domains or modalities often exhibit heterogeneity in node types, attribute formats, and structural patterns, which increases prompt length and introduces misalignment between textual and topological inputs [86, 220]. Existing methods struggle to handle these inconsistencies, as they typically rely on shallow normalization or isolated attribute encoding, and LLMs themselves lack permutation invariance when processing graph structures [252]. To address these limitations, future efforts should explore context-aware modeling that jointly encodes multimodal attributes, local connectivity, and task-specific signals. Symmetry-preserving techniques such as contrastive learning or explicit prompt designs can help models distinguish meaningful structural variations from order artifacts [256]. Additionally, tighter LLM–GNN integration—via dual encoders or knowledge distillation from graph-level GNNs—may provide a more balanced trade-off between semantic richness and structural fidelity [261, 268].
4.1.4 Dynamic Instability
In dynamic graphs, frequent structural or attribute updates introduce substantial overhead for model recomputation and pose challenges for maintaining temporal consistency [314]. Current LLM-based approaches often rely on full re-encoding or static snapshots, which fail to reflect evolving semantics and result in stale or inaccurate reasoning [322]. To overcome these limitations, future research should incorporate time-aware prompting, localized subgraph updates, and selective verification to reduce redundant computation [368]. Caching frequently accessed substructures or intermediate representations can further improve efficiency. More importantly, integrating GNNs with LLMs for temporal causal reasoning, along with pretraining strategies that embed temporal priors, may help models adapt to semantic drift and support robust inference over evolving graph states [321, 318].
4.2 Explainability, Fairness, and Faithfulness
LLM-based graph systems pose serious challenges in explainability, fairness, and faithful reasoning—especially in scenarios involving incomplete data, long-tail distributions, domain shifts, and temporal updates. While current approaches offer partial solutions, their limitations in interpretability, causal alignment, and robustness across scenarios call for more principled advancements.
4.2.1 Incompleteness
Incomplete graphs often result in LLM outputs that lack faithfulness, as the underlying reasoning is not grounded in observable evidence [134, 135, 121]. This issue is exacerbated by decoding processes that fail to enforce topological constraints and by the absence of systematic post-generation verification. Future work should incorporate structure-constrained decoding objectives and apply graph consistency losses during training [136], while complementing generation with post-hoc validation using GNN-based scorers or confidence filters [91, 93]. Furthermore, reasoning fidelity deteriorates significantly when only partial graph observations are available [135]. To address this, prompting strategies should explicitly encode subgraph structures to help LLMs reason over local topological contexts. In addition, iterative refinement mechanisms that alternate between retrieval, reasoning, and validation can simulate GNN-style aggregation while supporting auditable evidence chains.
4.2.2 Imbalance
Graph imbalance not only degrades predictive performance but also undermines explanation stability and reliability. Due to sparse semantics and weak connectivity, LLMs may struggle to generate faithful rationales for long-tail nodes. Moreover, existing interpretability tools rarely account for such imbalance, leading to inconsistent explanations under input perturbations. Future work should explore causal mechanisms underlying topological sparsity and use LLMs to generate logically consistent virtual subgraphs to restore class-level balance [187]. At the same time, interpretability evaluation should emphasize stability: explanations for minority-class nodes should remain consistent across counterfactual edits or input augmentations, and validation modules should reject rationales that do not correspond to truly discriminative substructures.
4.2.3 Cross-Domain Heterogeneity
Inconsistencies in label semantics, graph patterns, and textual attributes across domains make it difficult for LLMs to produce generalizable explanations. These mismatches often introduce domain-specific biases that harm fairness and reasoning quality [371, 372]. Current models lack the ability to maintain explanation coherence when transferred across domains [373, 374]. Future systems should develop domain-invariant rationale templates and establish multi-level auditing frameworks covering embedding spaces, output distributions, and textual justifications to detect and mitigate explanation drift. Additionally, aligning the structural reasoning paths produced by GNNs with the textual rationales generated by LLMs can bridge the gap between semantics and topology and enhance explanation faithfulness.
4.2.4 Dynamic Instability
The evolving nature of real-world graphs introduces temporal variations that challenge the stability and trustworthiness of model explanations. As nodes and edges change over time, previously valid rationales may become outdated or misleading, yet current systems lack temporal auditing mechanisms. Future work should introduce rolling, time-aware auditing protocols that explicitly link model outputs with graph update events [371]. Moreover, enabling executable explanations—where LLM-generated rationales are translated into symbolic constraints or verifiable subgraph traces—can improve the reliability and traceability of reasoning under temporal drift [375].
4.3 Security, Robustness, and Governance
As graph–LLM systems become increasingly applied in high-stakes domains such as finance, healthcare, and recommendation, they face growing vulnerabilities arising from incomplete information, class imbalance, cross-domain inconsistency, and evolving graph structures. These challenges expose the system to risks including data poisoning, prompt injection, structure manipulation, and adversarial generalization, many of which are amplified by the flexible yet opaque behavior of LLMs.
4.3.1 Incompleteness
In settings where graph data is sparse or noisy, even minor perturbations to node or edge attributes can lead to significant shifts in completion outcomes, compromising the integrity of the generated graph [376, 377]. Similarly, LLM-augmented retrieval systems are vulnerable to prompt-level attacks, where indirect injection or document poisoning misleads the model’s response [378]. To improve robustness, future work should investigate multi-layer defenses, including data-level sanitization and edge pruning, model-level protection through graph denoising techniques such as GNNGuard [379], and certification-based methods like randomized smoothing to quantify the system’s tolerance under adversarial conditions [380, 381].
4.3.2 Imbalance
Nodes associated with rare categories or low connectivity often lack sufficient structural context, making them particularly susceptible to adversarial influence; subtle perturbations can alter neighborhood aggregation or trigger class boundary shifts [382]. Moreover, few-shot trigger constructions may induce undesirable distributional shifts in LLM outputs. To address these issues, future systems should adopt targeted defenses for long-tail nodes, including structural consistency checks and trigger detection via adversarial training regimes that expose the model to rare-case manipulations.
4.3.3 Cross-Domain Heterogeneity
When LLM-based models are deployed across domains with differing label semantics, structure formats, or template designs, inconsistencies can be exploited to induce behavior shifts or prompt corruption. Existing systems often lack robust controls over prompt templates or graph encodings, making them vulnerable to injection and escalation attacks [378]. Looking forward, stronger domain-level governance is needed, including adversarial evaluation prior to deployment, enforcement of signed prompt templates, and whitelist-based control over structural tokens to prevent unauthorized modifications.
4.3.4 Dynamic Instability
Dynamic graphs introduce further complexity, as attackers may inject or remove connections during critical time windows or log adversarial traces to shape long-term model behavior [382]. These time-sensitive attacks are often stealthy and hard to reverse, leading to compounding damage [381]. Future directions should prioritize rolling-window detection and temporal rollback mechanisms, coupled with causal tracing of performance degradation to structural changes. Additional safeguards such as time-aware scoring and path-verifiable explanation generation can help separate legitimate evolution from malicious interference.
@@ 5 Conclusion
This survey has presented a comprehensive and systematic review of how recent advances in LLMs can be leveraged to address four fundamental, data-centric challenges in graph learning: incompleteness of structures or attributes, severe imbalance in node and edge distributions, cross-domain heterogeneity in semantics and structure, and dynamic instability arising from evolving topologies and interactions. To achieve this, we conducted an extensive literature collection and categorization, and organized representative methods. We further summarized benchmark datasets and evaluation metrics used in existing studies, assessed empirical trends. In addition, we identified open technical challenges and outlined promising future research directions.
We have synthesized a broad spectrum of LLM–graph integration strategies under this framework, demonstrating how LLMs bring distinctive capabilities that complement purely graph-based approaches. In the context of incompleteness, LLMs apply semantic reasoning and draw upon external knowledge to infer missing attributes and relationships, serving as intelligent imputers. For imbalanced graphs, they can generate synthetic samples and enrich feature spaces, thereby enhancing minority-class representations and mitigating bias. In heterogeneous graph scenarios, LLMs facilitate the unification of disparate modalities and domain-specific schemas into coherent embeddings, enabling effective cross-domain alignment. For dynamic graphs, their contextual and temporal reasoning allows for anticipating structural changes, explaining evolution, and supporting continuous adaptation.
@@
Looking ahead, the integration of LLMs with graph learning still faces important open questions. Improving efficiency and scalability is essential for large-scale or real-time applications, while enhancing interpretability and trustworthiness will be critical for deployment in high-stakes domains such as healthcare and finance. Developing mechanisms for continual adaptation without catastrophic forgetting remains a significant challenge, as does bridging the gap between textual knowledge encoded in LLMs and structural graph signals. In addition, controlling hallucination in LLM-generated graph content is vital to ensure both semantic validity and structural consistency.
@@
In conclusion, the convergence of LLMs and graph learning marks a promising new direction, combining the deep semantic understanding of natural language processing with the structured relational modeling of graph machine learning. This synergy has already yielded models that are more robust, knowledgeable, and adaptable than those based on either technology alone. As research advances, we anticipate rapid progress toward graph learning systems with greater generality and intelligence, capable of reasoning effectively over the rich, dynamic networks that underpin real-world data.
@@ Acknowledgements
This work was supported in part by the Shenzhen Science and Technology Program under Grant JCYJ20240813151129038; in part by the National Natural Science Foundation of China under Grant 52172350, Grant 51775565, and Grant W2421069; in part by the Guangdong Basic and Applied Research Foundation under Grant 2022B1515120072; and in part by the Science and Technology Planning Project of Guangdong Province under Grant 2023B1212060029.